Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Entry   Shift summary, posted by Giuseppe Mercurio, Diego, Martin, Amina on 05 May 2023, 06:33 
    Reply   Re: Shift summary, posted by Giuseppe Mercurio on 05 May 2023, 07:23 optical_sync_2023-05-05_07-21-35.png
Message ID: 179     Entry time: 05 May 2023, 06:33     Reply to this: 180
Author: Giuseppe Mercurio, Diego, Martin, Amina 
Group: SCS 
Subject: Shift summary 

achievements:

  • inserted the magnet and measured delay scan at the maximum magnetic field with opposite directions
  • data analysis showed that runs measured during the afternoon/evening led to sample degradation (the sample shows a dark spot, change in the scattering signal is also cearly visible)
  • corrected spatial overlap
  • moved to gresh spot and did delay scan at lower OL fluence (Tr = 35% instead of 70%)

problems:

  • shortly before 23:00 the optical laser synchronization was lost for a bit more than 2 hours, then it came back shortly after 1 am at that time the PP laser phase shifter when optically lock moved from 1018 to 229.8 apparently on its own, and it was realized few hours later. We moved it back to the original value, but this was not sufficient, we had to move by several ns and we were not able to move where we wanted, Laser OCD was called a7 7:35.
ELOG V3.1.4-7c3fd00